Only 4 of us today but a productive day. Graham and Richard spent time playing with the new Induction Heater for soldering the ends onto the power and traction cables for D9009. After a good few trials and alterations to the jig they have made they eventually were getting the required results. The key criteria are to get the solder hot enough to get it right down to the bottom of the brass fitting but then not to disturb it for a good while to allow it to set. If it moves and causes a crack below the surface this can cause bad conduction or even failure.
Bill continued with refitting the terminal boxes onto the reworked Cast Bogies for 9. Geoff has cleaned, painted a varnished them and their internal components with Anti tracking Varnish, Richard gave some assistance at the start of the day whilst Graham set up the Induction Heater. Once Bill had completed this task, he moved on to complete the cleaning, surface dust, from the bogies which Steve P started last week, looking excellent now. These will be installed under 9 over the next few weeks.
Meanwhile Mike continued with some of Darrell's Depot Tasks by consolidating the contents of a few pallets and putting them back into the racking with the least required items near the top row, several more identified. The PPE cabinet was relocated to a less obtrusive place. Following the PU swap in 2 weeks' time the spare refurbished generator will be placed on the splitting stand, this will create even more working space. The 2 forklift trucks have now been given new homes to free up more space. Just prior to lunch Bill and Richard broke off from what they were doing to 'do their bit' towards the external tidy and the remainder of the scrap metal was put in the skip and the scrap wood was sorted and tidied ready for breaking up and used in the new Depot Heaters.